A woman earns $1,350 in interest from two accounts in a year. If she has three times as much invested at 7% as she does at 6%, how much does she have in each account?

Answer by mananth(16946)

6.00% ------------ 1 x
7.00% ----------- 3 x
Interest = 1350
Interest equation

6.00% * 1.00 x + 7.00% * 3 x = 1350
0.060 x + 0.210 x = 1350
0.27 x= 1350
x= 1350 / 0.27
x= 5000

3 x = 15000

CHECK
6.00% x= 0.06 * 5000 = 300
7.00% ----------- 0.07 * 15000 = 1050

m.ananth@hotmail.ca